Bob is 9 years old and his brother is five. How many years will go by before bob is five-fourths his brother's age?
Answer by fractalier(6550)

Call the number of years in the future, x.
Then we can write
9 + x = (5/4)(x + 5)
9 + x = (5/4)x + 25/4
Now subtract x and subtract 25/4 from 9 and get
11/4 = 1/4 x
x = 11 years from now